1 1 1
- = - + -
y 56.1 15.3

Answer by stanbon(75887)

1 1 1
- = - + -
y 56.1 15.3
-----------------------------
(1/y) = (15.3+56.1)/(56.1*15.3)
---
1/y = 71.4/(858.330)
y = (858.33/71.4)
y = 12.0214
================
Cheers,
Stan H.